# SliM-LLM: Salience-Driven Mixed-Precision Quantization for Large Language Models

### v1

@@ 4 Experiments
4.2 Ablation Results.
We conduct a detailed ablation study to illustrate the benefits of bit-width allocation and the impact of each component. Fig. 5(a) compares three strategies for allocating bit-widths across groups, including random allocation, head-tail allocation by spatial order, and our proposed SBA. When the average bit-width remains constant, random and head-tail mixed-precision allocation prove ineffective and even result in performance degradation, as shown in Fig. 5(a). In contrast, SBA consistently delivers significant improvements in post-quantization performance, validating the efficacy of our mixed-precision approach. Fig. 5(b) presents the ablation effects of SBA and SQC, demonstrating that both methods, based on the perception of global and local salience, enhance quantization performance. SBA is particularly effective in smaller models, and combining these two methods can further boost capabilities of LLMs. We also provide the detailed ablation results on group size in Appendix E.
Figure 5: Ablation results on OPT models. Random means randomly selecting the same number of lower/higher-bit groups; head-tail denotes using the head groups as the lower-bit and the same number of tails as the higher-bit on the original sequence of group.
@@
FP16 - 12.6G 14.4G 5.68 69.2 24.3G 27.1G 5.09 52.5 12.7G 14.6G 5.47 69.3
3-bit GPTQ 3.2G 5.1G 6.55 83.4 5.8G 8.7G 5.62 57.6 3.2G 5.2G 6.29 56.3
SliM-LLM 3.2G 5.2G 6.40 79.1 5.8G 8.8G 5.48 48.5 3.2G 5.4G 6.26 55.9
2-bit GPTQ 2.2G 4.1G 152.31 83.9 4.0G 7.5G 20.44 92.6 2.2G 4.1G 60.45 83.6
SliM-LLM 2.3G 4.4G 14.58 61.2 4.1G 7.8G 8.87 73.7 2.3G 4.1G 16.01 64.4
@@ 4.3 Efficient Inference on Device
We utilize the open-source AutoGPTQ to extend CUDA kernel supporting experimental mixed-precision inference, with detailed process in Appendix B.2. We evaluate the deployment performance of LLaMA-7/13B and LLaMA-2-7B under 2/3-bit settings. The results indicate that our mixed-precision approach maintains a good compression rate on GPUs and significantly enhances model accuracy, only with a slight decrease in inference speed on the A800 (due to the inference alignment of different bit-width). Since current 1-bit operations lack well hardware support, additional consumption of storage and computation is required on device. There remains considerable scope for optimization in mixed-precision computing, and we aim to further improve this in future work.
@@ 5 Conclusion
In this work, we introduce SliM-LLM, a structured mixed-precision PTQ framework tailored for LLMs, designed to enhance performance with low-bit weights in a deployment-friendly manner. The essence of SliM-LLM lies in employing the Salience-Determined Bit Allocation to dynamically allocate bit widths, thereby improving the preservation of global salience information. Within groups, the Salience-Weighted Quantizer Calibration is designed to enhance local information perception, further minimizing the loss associated with locally salient weights. Experiments validate the effectiveness of SliM-LLM, showing notable accuracy improvements across various LLMs, and ensuring efficiency in inference. In conclusion, SliM-LLM is versatile and can be seamlessly integrated with different quantization frameworks and successfully improves the performance of LLMs supporting practical deployment in resource-constrained environments.
Appendix A Broader Impacts and Limitations
A.1 Broader Impacts
@@ A.2 Limitations
A.3 Experiments Reproducibility
Our code is included in the supplementary materials. For instructions on how to reproduce various experiments, please refer to the accompanying code scripts and algorithm description in our paper. We also provide the download and use details of the datasets mentioned in the experiment part.
@@ B.1 Detailed Implementation
In this section, we present the specific implementation details of SliM-LLM, which utilizes GPTQ [16] as its backbone for mixed-precision quantization and incorporates both SBA and SQC. SliM-LLM+ is consistent with SliM-LLM in SBA computations but does not include the SQC component, instead retaining learnable weight clipping (LWC) approach in OmniQuant [38] for gradient optimization.
@@
2:  $e\coloneqq\operatorname{inf}$ //  bit-width searching error
@@
8:  for $\gamma\in~{}[1-\lambda,1+\lambda]$ with $2n$ slices do
9:     $\Delta\coloneqq\gamma(w_{\mathrm{max}}-w_{\mathrm{min}})/(2^{g_{s}}-1)$
10:     $z\coloneqq-\lfloor(\gamma w_{\mathrm{min}})/{\Delta}\rceil$
@@ B.2 Mixed Bit Storage and Computing
We developed a framework for storage and inference deployment supporting mixed-precision quantization based on AutoGPTQ. The deployment process is as follows. After completing mixed-precision quantization with SliM-LLM, it outputs scales, zeros, and group-wise bit-width generated during the quantization process to identify the quantization parameters and precision of each group in the Linear Projection weights. AutoGPTQ then packs the weights and zeros into integer-compressed representations (denoted by $\hat{\boldsymbol{w}}_{\mathrm{int}}$ and $\hat{\boldsymbol{z}}_{\mathrm{int}}$ respectively) based on the precision of different groups, significantly reducing storage and operational bit-width. After the quantized weights are packed, AutoGPTQ loads the model onto the GPU, where the mixed precision quantization kernel on the GPU performs dequantization on the weights and zeros of different groups and calculation with input activation, ultimately producing the final output.
In the mixed-precision deployment of AutoGPTQ, the weight memory layout is organized by group, with each group sharing the same precision, which is shown in Fig. 6. Within each group, elements with the same precision are packed as integers, eliminating the need for additional padding, which saves space. Given that the bit-width of integers is a power of 2, this is compatible with group size that is also a power of 2. For instance, even with the odd-bit such as 3-bit storage, integers can store these numbers without padding, as the commonly used group size is 128, a multiple of almost all definition of integer type. This ensures that elements within a group fully utilize the space provided by integers, without storing numbers of different precision within the same integer. $\hat{\boldsymbol{z}}_{\mathrm{int}}$ follow the original logic of AutoGPTQ but are packed with a uniform precision along the channel direction for ease of use. Other tensors, like scales, remain in the same floating-point format to ensure the correctness of dequantization calculations.
To indicate the precision of each group, we also introduce an additional array to store bit-width of each group, where each number is represented as a 2-bit value aggregated into integers, marking the quantization precision of each group for accurate reconstruction. We use cumulative calculations to determine the starting index of each group, ensuring correctness despite changes in $\hat{\boldsymbol{w}}_{\mathrm{int}}$ height and starting indices caused by varying precision. Using the above methods to store the quantized weights, zeros, and additional bit arrays effectively reduces memory usage during model storage and loading, thereby lowering the resource overhead required for model deployment.
Once the weights are packed, we follow the modified AutoGPTQ logic for GPU inference. The GPU processes and dequantizes the weights group by group for computation. During GPU computation, a thread dequantizes a segment of continuous memory data in one column of $\hat{\boldsymbol{w}}_{\mathrm{int}}$ and performs vector dot product calculations with the input activation shared within the block, accumulating the results in the corresponding result matrix. When threads form a logical block, the block handles the computation and reduction of a continuous channel region. We complete the linear layer computation by iterating through all logical blocks. Leveraging AutoGPTQ’s initial logic and CUDA Warp’s 32-thread units, we ensure similar code structure and data access logic for threads within each warp when group size is 128. This method was primarily conducted to validate feasibility os SliM-LLM, demonstrating that the mixed precision quantization with integer packing does not cause additional computational overhead, indicating the efficiency and accuracy advantage of SliM-LLM. In summary, by dividing weight into several structured precision blocks and employing a reasonable GPU utilization strategy, Slim-LLM balances performance and efficiency.
@@ Appendix C Searching Details of Group-Wise Salience-Determined Bit Allocation
We optimize the mixed-precision configuration based on the output information entropy (KL-divergence), searching for the optimal compensation bit-width ratio as shown in Eq. (4).
Initially, we rank each group by their average salience, a metric for quantization, and employ a double-pointer that moves simultaneously from both the beginning (lowest salience) and end (highest salience) of the sorted list. This ensures an equal number of groups at low and high bit-widths, effectively balancing the global average bit-width compensation. We then calculate the relative entropy under the corresponding precision ratio and search for the optimal ratio. Fig 7 displays the search error curves related to the $2^{nd}$, $10^{th}$, and $15^{th}$ Transformer layers in the OPT1.3B model, showcasing the search curves for certain self-attention layers (Query, Key, Value, FC2).
@@
Appendix D Extension Ablation on SQC
In this section, we visualize the effectiveness of SQC in mitigating the degradation of information in locally salient weights. We observed the absolute error of weights in a randomly selected channel of the quantized OPT-1.3B model. As shown in Fig. 8, the overall absolute error of the weights post-quantization with a standard quantizer was 0.0055, while with SQC it was reduced to 0.0039. This further demonstrates that the search parameter $\gamma$, as applied in Eq. (5), effectively optimizes the quantizer parameters, thereby reducing quantization errors.
@@ Appendix D Extension Ablation on SQC
Figure 8: Absolute channel error of the weight of the OPT-1.3B model. The red line represents the quantization error for the locally salient weights, and the gray represents other weights. (a) Vanilla quantizer error on the $794^{th}$ channel of OPT-1.3B. (b) SQC error on the $794^{th}$ channel of OPT-1.3B
Appendix E Extension Ablation on Quantization Group-Size
@@
Appendix F Extension on Salience Channel Clustering
F.1 Discussion of Theorem 1
@@
Given the input calibration activation $\boldsymbol{x}\in\mathbb{R}^{t\times m}$ with an outlier value $x_{p,q}^{*}\gg x_{i,j},\forall i\in[0,t],i\neq p,\forall j\in[0,m],j\neq q$ at the position of token-$p$ and channel-$q$. The diagonal elements of $\boldsymbol{x}\boldsymbol{x}^{\top}$ also shows outlier value at $[\boldsymbol{x}\boldsymbol{x}^{\top}]_{p,p}$, as ${x_{p,q}^{*}}^{2}$ produced by $[\boldsymbol{x}\boldsymbol{x}^{\top}]_{i,i}=\sum_{k=1}^{n}x_{i,k}^{2}$ only appearing at position $(p,p)$, which further leads to the parameter salience larger at the $p^{th}$ channel of weight, where $\delta_{:,p}>\delta_{:,k},\delta_{:,k}=\frac{w_{:,k}^{2}}{[\boldsymbol{H}^{-1}]_{k,k}^{2}},\forall k\in[0,t],k\neq p$.
@@
Given $\boldsymbol{x}\in\mathbb{R}^{t\times m}$ with outlier value $x_{p,q}^{*}$ at token-$p$ and channel-$q$, $p\in[0,t]$ and $q\in[0,m]$, and other elements with small magnitude $x_{i,j}$, where $x_{p,q}^{*}\gg x_{i,j}$ and $i,j\neq p,q$. We can get the Hessian matrix with Levenberg-Marquardt [29] approximation in Eq. (3):
$\begin{pmatrix}x_{11}&x_{12}&x_{13}&\cdots&x_{1m}\\ x_{21}&x_{22}&x_{23}&\cdots&x_{2m}\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ \vdots&\vdots&\vdots&\boldsymbol{x_{p,q}^{*}}&\vdots\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ x_{t1}&x_{t2}&x_{t3}&\cdots&x_{tm}\\ \end{pmatrix}\cdot\begin{pmatrix}x_{11}&x_{12}&\cdots&x_{1t}\\ x_{21}&x_{22}&\cdots&x_{2t}\\ \vdots&\vdots&\ddots&\vdots\\ \vdots&\vdots&\boldsymbol{x_{p,q}^{*}}&\vdots\\ \vdots&\vdots&\ddots&\vdots\\ x_{m1}&x_{m2}&\cdots&x_{mt}\\ \end{pmatrix}=\begin{pmatrix}x_{11}^{2}..&\cdots&\cdots&\cdots\\ \vdots&\ddots&\cdots&\vdots\\ \vdots&\vdots&\boldsymbol{x_{p,q}^{*}}^{2}..&\vdots\\ \cdots&\cdots&\cdots&\ddots\\ \end{pmatrix}$ (7)
where ${x_{p,q}^{*}}^{2}$ only appears at position $\boldsymbol{H}_{p,p}$. And following SparseGPT [15], the inverse matrix of $\boldsymbol{H}$ can be formulated as:
$\delta_{i,j}=\frac{w_{i,j}^{2}}{[\operatorname{diag}((\boldsymbol{x}\boldsymbol{x}^{\top}+\lambda\boldsymbol{I})^{-1})]^{2}}$ (8)
where $(\boldsymbol{x}\boldsymbol{x}^{\top}+\lambda\boldsymbol{I})^{-1}$ is the new representation of Hessian matrix $\boldsymbol{H}$ for the layer-wise reconstruction problem, and $\lambda$ is the dampening factor for the Hessian to prevent the collapse of the inverse computation. Additionally, in accordance with the configuration in LLMs [15, 16, 39], the value of $\lambda$ set is extremely small ($\lambda\leq e^{-1}$), while the values located at the diagonal of Hessian are large. Therefore, only considering the influence of diagonal elements [39], we can further approximate salience as:
$\delta_{i,j}=\frac{w_{i,j}^{2}}{[\operatorname{diag}((\boldsymbol{x}\boldsymbol{x}^{\top}+\lambda\boldsymbol{I})^{-1})]^{2}}\approx\frac{w_{i,j}^{2}}{[(\operatorname{diag}(\boldsymbol{x}\boldsymbol{x}^{\top}))^{-1}]^{2}}=(w_{i,j}\cdot||\boldsymbol{x}_{j}||_{2}^{2})^{2}$ (9)
Here the diagonal of $\boldsymbol{x}\boldsymbol{x}^{\top}$ is $\operatorname{diag}(||\boldsymbol{x}_{j}||_{2}^{2})$, and $||\boldsymbol{x}_{j}||_{2}$ evaluates the $\ell_{2}$ norm of $j^{th}$ channel across different tokens. Consequently, it can be summarized that when there is an outlier value at the position of the token-$p$ and channel-$q$, $||\boldsymbol{x}_{p}||_{2}$ is primarily influenced by $x_{p,q}^{*}$. Additionally, since the activation values are relatively large and the differences in weight values are comparatively small, the $p^{th}$ channel of weights will also exhibit salience. ∎
F.2 Distribution of salience, activation and weight magnitude

### v2

@@ 4.1 Main Results
Figure 5: Ablation results on OPT models. Random refers to randomly selecting lower- and higher-bit groups, while head-tail assigns lower-bit precision to the head groups and higher-bit precision to an equal number of tail groups in the original sequence.
4.2 Ablation Results
Table 4: WikiText2$\downarrow$ performance of SBA and ILP on LLaMA.
Method #W 7B 13B 30B 65B
ILP 2-bit 17.55 9.51 9.27 7.46
SBA 2-bit 14.58 8.87 7.33 5.90
Abliation of SBA and SQC. We conduct a detailed ablation study to illustrate the benefits of bit-widths allocation and the impact of each component. Fig. 5(a) compares three strategies for allocating bit-widths across groups, including random allocation, head-tail allocation by spatial order, and our proposed SBA. When the average bit-widths remains constant, random and head-tail mixed-precision allocation prove ineffective and even result in performance degradation, as shown in Fig. 5(a). In contrast, SBA consistently delivers significant improvements in post-quantization performance, validating the efficacy of our mixed-precision approach. Fig. 5(b) presents the ablation effects of SBA and SQC, demonstrating that both methods, based on the perception of global and local salience, enhance quantization performance.
Compare of SBA and ILP. We compare the performance between the ILP model in HAWQ v2 (Dong et al., 2019) and SBA on the LLaMA model. Tab. 4 shows that SBA achieves comprehensive performance superiority on LLaMA. We observed that under a 2-bit scenario, ILP ensures an equal number of 1-bit and 3-bit groups within the search space {1-bit, 2-bit, 3-bit}. The advantage of ILP lies in a broader selection range for target bit-widths, but under commonly used fixed integer bit-widths (e.g. 2-bit, 3-bit), SBA’s double-pointer search strategy based on output feature KL proposed by SBA can achieve a more optimal matching strategy.
@@ 4.2 Ablation Results
FP16 - 12.6G 14.4G 5.68 69.2 24.3G 27.1G 5.09 52.5
3-bit GPTQ 3.2G 5.1G 6.55 83.4 5.8G 8.7G 5.62 57.6
SliM-LLM 3.2G 5.2G 6.40 79.1 5.8G 8.8G 5.48 48.5
2-bit GPTQ 2.2G 4.4G 152.31 83.9 4.0G 7.6G 20.44 92.6
SliM-LLM 2.3G 4.4G 14.58 61.2 4.1G 7.8G 8.87 73.7
@@ 4.3 Efficient Inference on Device
We utilize the open-source AutoGPTQ to extend CUDA kernel supporting experimental mixed-precision inference, with detailed process in Appendix B.2. As shown in Tab. 5, we evaluate the deployment performance of LLaMA-7/13B and LLaMA-2-7B under 2/3-bit settings. The results indicate that our mixed-precision approach maintains a good compression rate on GPUs and significantly enhances model accuracy, only with a slight decrease in inference speed on the A800. We also deploy the LLaMA-2-70B in Tab. 14. Since current 1-bit operations lack well hardware support, additional consumption of storage and computation is required on device. There remains considerable scope for optimization in mixed-precision computing, and we aim to further improve this in future work.
@@ 5 Conclusion
In this work, we propose SliM-LLM, a group-wise mixed-precision PTQ framework for LLMs, designed to optimize performance with low-bit weights while maintaining deployment efficiency. The core of SliM-LLM is the Salience-Determined Bit Allocation, which dynamically assigns bit widths to preserve global salience information. Additionally, the Salience-Weighted Quantizer Calibration enhances local information retention, reducing the impact of quantization on locally salient weights. Experimental results demonstrate that SliM-LLM significantly improves accuracy across various LLMs while ensuring inference efficiency. Overall, SliM-LLM is a versatile solution that integrates seamlessly with existing quantization frameworks, enabling practical deployment in resource-constrained environments.
Acknowledgments
The authors acknowledge the Innovation and Technology Fund from Innovation and Technology Commission of Hong Kong SAR Government (Mainland-Hong Kong Joint Funding Scheme MHP/053/21), Shenzhen-Hong Kong-Macau Technology Research Programme (SGDX20210823103537034) from Shenzhen Science and Technology Innovation Commission, and Seed Funding for Strategic Interdisciplinary Research Scheme from the University of Hong Kong (HKU) for supporting this work. This work has also been supported by Hong Kong Research Grant Council - Early Career Scheme (Grant No. 27209621), General Research Fund Scheme (Grant No. 17202422, 17212923), the Swiss National Science Foundation (SNSF) project 200021E_219943 Neuromorphic Attention Models for Event Data (NAMED)
Impact Statement
@@ B.1 Detailed Implementation
In this section, we present the specific implementation details of SliM-LLM, which utilizes GPTQ (Frantar et al., 2022) as its backbone for mixed-precision quantization and incorporates both SBA and SQC. SliM-LLM+ is consistent with SliM-LLM in SBA computations but does not include the SQC component, instead retaining learnable weight clipping (LWC) approach in OmniQuant (Shao et al., 2023) for gradient optimization.
Algorithm 1 Main Framework of SliM-LLM.
Input: $\boldsymbol{w}\in\mathbb{R}^{n\times m}$ FP16 weight
          $\boldsymbol{x}_{F}\in\mathbb{R}^{t\times m}$ calibration data
          $\beta$ group size
          $\lambda$ hessian regularizer
          $N$ average bit-width
Output: $\hat{\boldsymbol{w}}_{q}$ quantized weight
3:  $\hat{\boldsymbol{w}}_{q}\coloneqq 0^{n\times m}$
5:  for $b=0,\beta,2\beta,...$ do
6:     $\boldsymbol{w}^{b}\coloneqq\boldsymbol{w}_{:,b:b+\beta}$
7:     $g_{b}\coloneqq\mathcal{G}[b]$
10:     GPTQ-error compensation:
13:  end for
14:  return $\hat{\boldsymbol{w}}_{q}$
@@
2:  $e\coloneqq\operatorname{inf}$ //  bit-widths searching error
@@
8:  for $\tau\in~{}[1-\lambda,1+\lambda]$ with $2n$ slices do
9:     $\Delta\coloneqq\tau(w_{\mathrm{max}}-w_{\mathrm{min}})/(2^{g_{s}}-1)$
10:     $z\coloneqq-\lfloor(\tau w_{\mathrm{min}})/{\Delta}\rceil$
@@ B.2 Mixed Bit Storage and Computing
We developed a framework for storage and inference deployment supporting mixed-precision quantization based on AutoGPTQ. The deployment process is as follows. After completing mixed-precision quantization with SliM-LLM, it outputs scales, zeros, and group-wise bit-widths generated during the quantization process to identify the quantization parameters and precision of each group in the Linear Projection weights. AutoGPTQ then packs the weights and zeros into integer-compressed representations (denoted by $\hat{\boldsymbol{w}}_{\mathrm{int}}$ and $\hat{\boldsymbol{z}}_{\mathrm{int}}$ respectively) based on the precision of different groups, significantly reducing storage and operational bit-width. After the quantized weights are packed, AutoGPTQ loads the model onto the GPU, where the mixed precision quantization kernel on the GPU performs dequantization on the weights and zeros of different groups and calculation with input activation, ultimately producing the final output.
@@
In the mixed-precision deployment of AutoGPTQ, the weight memory layout is organized by group, with each group sharing the same precision, which is shown in Fig. 6. Within each group, elements with the same precision are packed as integers, eliminating the need for additional padding, which saves space. Given that the bit-widths of integers is a power of 2, this is compatible with group size that is also a power of 2. For instance, even with the odd-bit such as 3-bit storage, integers can store these numbers without padding, as the commonly used group size is 128, a multiple of almost all definition of integer type. This ensures that elements within a group fully utilize the space provided by integers, without storing numbers of different precision within the same integer. $\hat{\boldsymbol{z}}_{\mathrm{int}}$ follow the original logic of AutoGPTQ but are packed with a uniform precision along the channel direction for ease of use. Other tensors, like scales, remain in the same floating-point format to ensure the correctness of dequantization calculations.
To indicate the precision of each group, we also introduce an additional array to store bit-widths of each group, where each number is represented as a 2-bit value aggregated into integers, marking the quantization precision of each group for accurate reconstruction. We use cumulative calculations to determine the starting index of each group, ensuring correctness despite changes in $\hat{\boldsymbol{w}}_{\mathrm{int}}$ height and starting indices caused by varying precision. Using the above methods to store the quantized weights, zeros, and additional bit arrays effectively reduces memory usage during model storage and loading, thereby lowering the resource overhead required for model deployment.
Once the weights are packed, we follow the modified AutoGPTQ logic for GPU inference. The GPU processes and dequantizes the weights group by group for computation. During GPU computation, a thread dequantizes a segment of continuous memory data in one column of $\hat{\boldsymbol{w}}_{\mathrm{int}}$ and performs vector dot product calculations with the input activation shared within the block, accumulating the results in the corresponding result matrix. When threads form a logical block, the block handles the computation and reduction of a continuous channel region. We complete the linear layer computation by iterating through all logical blocks. Leveraging AutoGPTQ’s initial logic and CUDA Warp’s 32-thread units, we ensure similar code structure and data access logic for threads within each warp when group size is 128. This method was primarily conducted to validate feasibility os SliM-LLM, demonstrating that the mixed precision quantization with integer packing does not cause additional computational overhead, indicating the efficiency and accuracy advantage of SliM-LLM. In summary, by dividing weight into several structured groups with mixed precision and employing a reasonable GPU utilization strategy, Slim-LLM balances performance and efficiency.
@@ Appendix C Searching Details of Group-Wise Salience-Determined Bit Allocation
We optimize the mixed-precision configuration based on the output information entropy (KL-divergence), searching for the optimal compensation bit-widths ratio as shown in Eq. (4).
Initially, we rank each group by their average salience, a metric for quantization, and employ a double-pointer that moves simultaneously from both the beginning (lowest salience) and end (highest salience) of the sorted list. This ensures an equal number of groups at low and high bit-widths, effectively balancing the global average bit-widths compensation. We then calculate the relative entropy under the corresponding precision ratio and search for the optimal ratio. Fig 7 displays the search error curves related to the $2^{nd}$, $10^{th}$, and $15^{th}$ Transformer layers in the OPT1.3B model, showcasing the search curves for certain self-attention layers (Query, Key, Value, FC2).
@@
Appendix D Evluatiion Function of SBA
In Tab. 6, we employ various objective functions and compare their performance in SBA across different models. Compared to the commonly used Mean Squared Error (MSE) loss, Kullback-Leibler (KL) divergence ensures the distribution of critical activation positions within the model from an information entropy perspective, making it a superior choice for the bit-widths allocation strategy in SBA for the OPT and LLaMA models. When computing KL divergence in this context, we first transform the layer outputs into probability distributions using softmax.
Table 6: Comparison of MSE and KL Divergence in SBA.
Method # W OPT-1.3B OPT-2.7B OPT-6.7B OPT-13B LLaMA-7B LLaMA2-7B
MSE 2-bit 32.50 27.58 15.14 13.28 21.94 16.86
KL Divergence 2-bit 30.71 13.26 11.27 10.12 14.58 16.01
Appendix E Extension Ablation on SQC
In this section, we visualize the effectiveness of SQC in mitigating the degradation of information in locally salient weights. We observed the absolute error of weights in a randomly selected channel of the quantized OPT-1.3B model. As shown in Fig. 8, the overall absolute error of the weights post-quantization with a standard quantizer was 0.0055, while with SQC it was reduced to 0.0039. This further demonstrates that the search parameter $\tau$, as applied in Eq. (5), effectively optimizes the quantizer parameters, thereby reducing quantization errors.
@@ Appendix E Extension Ablation on SQC
Figure 8: Absolute channel error of the weight of the OPT-1.3B model. The red line represents the quantization error for the locally salient weights, and the lightmauve represents other weights. (a) Vanilla quantizer error on the $794^{th}$ channel of OPT-1.3B. (b) SQC error on the $794^{th}$ channel of OPT-1.3B
@@
Appendix F Extension Ablation on Quantization Group-Size
Appendix G Extension on Salience Channel Clustering
G.1 Discussion of Theorem 1
@@
Given the input calibration activation $\boldsymbol{x}\in\mathbb{R}^{t\times m}$ with an outlier channel $\boldsymbol{x_{:,p}^{*}}\gg\boldsymbol{x}_{:,j},\forall j\in[0,m],j\neq p$ at the position of channel-$p$. The trace elements of $\boldsymbol{H}=\boldsymbol{x}^{\top}\boldsymbol{x}$ will show great outlier value at $(p,p)$, where $\boldsymbol{H}_{p,p}\gg\boldsymbol{H}_{j,j},\forall j\in[0,m],j\neq p$, as $\boldsymbol{H}_{p,p}$ is produced by $[\boldsymbol{x_{:,p}^{*\top}}\boldsymbol{x_{:,p}^{*}}]=\sum_{i=0}^{t}x_{i,p}^{*2}$, which further leads to the parameter salience larger at the $p^{th}$ channel of weight, where $\delta_{:,p}>\delta_{:,k},\delta_{:,k}=\frac{w_{:,k}^{2}}{[\boldsymbol{H}^{-1}]_{k,k}^{2}},\forall k\in[0,t],k\neq p$.
@@
Given $\boldsymbol{x}\in\mathbb{R}^{t\times m}$ with outlier channel $\boldsymbol{x_{:,p}^{*}}$, $p\in[0,m]$, and other elements with small magnitude $x_{i,j}$, where $x_{q,p}^{*}\gg x_{i,j}$ and $i,j\neq q,p$. We can get the Hessian matrix with Levenberg-Marquardt (Marquardt, 1963) approximation in Eq. (3):
$\boldsymbol{H}=\begin{pmatrix}x_{11}^{2}+..&\cdots&\cdots&\cdots\\ \vdots&\ddots&\cdots&\vdots\\ \vdots&\vdots&\boldsymbol{{x_{p,p}^{*}}^{2}}+..&\vdots\\ \cdots&\cdots&\cdots&\ddots\\ \end{pmatrix}$ (6)
where $[\boldsymbol{x_{:,p}^{*\top}}\boldsymbol{x_{:,p}^{*}}]$ will appears at position $\boldsymbol{H}_{p,p}$. And following SparseGPT (Frantar & Alistarh, 2023), the inverse matrix of $\boldsymbol{H}$ can be formulated as:
$\delta_{i,j}=\frac{w_{i,j}^{2}}{[\operatorname{diag}((\boldsymbol{x}^{\top}\boldsymbol{x}+\lambda\boldsymbol{I})^{-1})]^{2}}$ (7)
where $(\boldsymbol{x}^{\top}\boldsymbol{x}+\lambda\boldsymbol{I})^{-1}$ is the new representation of Hessian matrix $\boldsymbol{H}$ for the layer-wise reconstruction problem, and $\lambda$ is the dampening factor for the Hessian to prevent the collapse of the inverse computation. Additionally, in accordance with the configuration in LLMs (Frantar & Alistarh, 2023; Frantar et al., 2022; Sun et al., 2023), the value of $\lambda$ set is extremely small ($\lambda\leq\text{e}^{-1}$), while the values located at the diagonal of Hessian are large. Therefore, only considering the influence of diagonal elements (Sun et al., 2023), we can further approximate salience as:
$\delta_{i,j}=$ $\frac{w_{i,j}^{2}}{[\operatorname{diag}((\boldsymbol{x}^{\top}\boldsymbol{x}+\lambda\boldsymbol{I})^{-1})]^{2}}\approx$ (8)
$\frac{w_{i,j}^{2}}{[(\operatorname{diag}(\boldsymbol{x}^{\top}\boldsymbol{x}))^{-1}]^{2}}=(w_{i,j}\cdot||\boldsymbol{x}_{j}||_{2}^{2})^{2}$
Here the diagonal of $\boldsymbol{x}^{\top}\boldsymbol{x}$ is $\operatorname{diag}(||\boldsymbol{x}_{j}||_{2}^{2})$, and $||\boldsymbol{x}_{j}||_{2}$ evaluates the $\ell_{2}$ norm of $j^{th}$ channel across different tokens. Consequently, it can be summarized that when there is an outlier channel-$p$, the value of $||\boldsymbol{x}_{p}||_{2}$ is primarily influenced by $[\boldsymbol{x_{:,p}^{*\top}}\boldsymbol{x_{:,p}^{*}}]$. Additionally, since the activation values are relatively large and the differences in weight values are comparatively small, the $p^{th}$ channel of weights will also exhibit salience. ∎
G.2 Distribution of salience, activation and weight magnitude